OpenStudy (anonymous):

Hi. :) I was wondering how I would find the value of "x"? (Drawing the pic now.) ^__^

OpenStudy (anonymous):

..

OpenStudy (anonymous):



OpenStudy (anonymous):

I got 23/6. But maybe I did something wrong? O__O

jimthompson5910 (jim_thompson5910):

Assuming the endpoints of the inner segment lie on the midpoints of the triangle, this means that the length of the inner segment is exactly half of the base So (1/2)(8x) = 2x + 23 Now solve for x (1/2)(8x) = 2x + 23 4x = 2x + 23 4x-2x = 23 2x = 23 x = 23/2

OpenStudy (anonymous):

:D Ok, thank you Jim. ^__^

jimthompson5910 (jim_thompson5910):

yw